# CoDA: Instructive Chain-of-Domain Adaptation with Severity-Aware Visual Prompt Tuning

### v1

@@ 2.1 Semantic segmentation under adverse scenes within UDA
Since existing UDA works predominately focus on four adverse scenes (fog, rain, snow, and night), we divide them into three branches by the scene benchmarks that they evaluate. The first branch comprises methods focusing on parsing foggy scenes [38, 39, 7] trying to generate synthetic fog images to narrow the real-to-foggy gaps. Later works tried other paths like CuDA-Net [31] that reported disentangling gaps by introducing an intermediate domain, and FIFO [28] that presented an auxiliary network to help the segmentation model to learn fog-invariant features. Since the above works generalize well to rainy and snowy scenes, no methods are specializing in rainy and snowy scenes. Night-specialized methods [52] are contained in the second branch and they too focus on narrowing the gaps by introducing extra domain images like twilight images or extra information[8, 37, 40]. The final branch comprises models trying to solve multiple scenes including normal and adverse scenes. DAFormer [20] was the first to introduce the SegFormer-based [54] architecture to this field. Subsequently, some DAFormer-based methods [53, 4] emerged like STA [16] which achieves domain generalization in UDA, and HRDA [21] that proposes multi-scale features fusion network. Some works also take HRDA as the baseline model, such as MIC [22] using masks to help models learn context-level features. Different from previous methods, CoDA is the first UDA method to focus models on differences within adverse scenes to learn domain-invariant features.
2.2 CoT and its variants in unsupervised domain adaptation
@@
2.3 Visual prompt tuning in unsupervised domain adaptation
Prompt-based learning [29] initially is used to finetune large pre-trained models to downstream tasks in Natural Language Processing (NLP). Subsequently, the prompt has begun to be transferred to the field of Computer Vision (CV) bringing about a significant academic trend. Many methods [33, 26, 56, 59] first tried to mitigate prompt to Vision Language Models (VLM) in text form, and then VPT [25] first introduced "visual prompt" into CV as learnable vectors (soft prompt). [14] followed by other works [13, 12, 42] is the first work focusing on domain adaptation and employing prompt. Most relevant to our work are the methods that focus on UDA or DG for adverse scene understanding, such as [11] employing text features to strengthen features achieving a novel style augmentation implicitly, and [46] combining clip [33] to conduct semantic augmentation. Our Meta-Visual Prompts with Meta-Adapters are explicitly added to images and guided by Severity to instruct models to learn domain-invariant features and ignore the scene-specific features.
@@ 3.3 Instructive Chain-of-Domain Adaptation
The first key point in scene-level CoD aims to simply divide all adverse scenes into Easy and Hard two-level scenes to construct two intermediate domains, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Take Cityscapes to ACDC experiment as an example, ACDC represents the target domain $T$ consists of fog, rain, snow, and night scenes, so we directly choose daytime scene images (fog, rain, snow, and $X_{G}$) as easy scene images to construct $X_{\mathcal{M}_{1}}$ and use night images to build $X_{\mathcal{M}_{2}}$. The data composition of Cityscapes to ACDC can be represented as Fig 3 (a). Thus, the CoD adaptation process transforms from $S\rightarrow T$ into $S\rightarrow\mathcal{M}_{1}\rightarrow\mathcal{M}_{2}\rightarrow T^{\prime}$.
@@ 4.1 Datasets and Benchmarks
Cityscapes (CS) [6] is a real-world dataset captured with normal driving scenes in 50 urban areas and contains 2,975 for training, 500 for validation, and 1,525 for testing. Foggy Zurich (FZ) [38] and Foggy Driving (FD) [39] are two foggy real-world datasets. FZ contains 3808 images with light and medium fog. Notably, only 40 images are for testing in FZ. Unlike FZ, FD is a dataset purely for testing as it contains 101 labeled images. Dark Zurich (DZ) [37] used for performance evaluation under night scenes comprises 8,779 images captured during nighttime, twilight, and daytime including 50 validation images and 151 test images. Nighttime Driving (ND) [8] contains 50 nighttime images with coarsely annotated ground truth. BDD100K-Night (BD) [57] is a subset of the BDD100K, we only use 87 test images. Similar to FD and BD, ND is also used for testing. ACDC [41] is a dataset with testing benchmarks under four adverse scenes: fog, rain, snow, and night. There are 400 training images, 100 validation images (including 106 night images), and 500 testing images in every scene. Besides, ACDC also contains 1600 clean images called ACDC-ref.
@@
Models Adverse Scenes (mIoU)
@@
SFSU [39] - 35.7 46.3 - - - - - - -
CMAda3 [7] - 46.8 49.8 - - - - - - -
CuDA-Net [31] 55.6 49.1 53.5 - - - - - - -
FIFO [28] - 48.4 50.7 - - - - - - -
FogAdapt [23] - 50.6 53.4 - - - - - - -
SAM-EDA [50] - - 56.4 - - - - - - -
GCMA [37] - - - - 42.0 45.6 33.2 - - -
MCGDA [40] - - - - 42.5 49.4 34.9 - - -
DANNet [52] - - - - 44.3 47.7 - - - 50.0
@@
AdaptSeg [44] - 26.1 37.6 - 30.4 34.5 22.0 - - -
DAFormer [20] 48.9 40.8 - 44.7 48.5 51.8 33.9 59.9 53.7 55.4
SePiCo [53] 58.5 - - 50.5 54.2 56.9 40.6 66.1 57.9 59.1
STA [16] 60.2 46.9 54.9 48.4 - - - 61.3 58.0 60.9
HRDA [21] 69.9 46.0 - 53.1 55.9 - - 73.6 69.5 68.0
MIC [22](baseline) 67.0 53.3 56.6 57.2 60.2 58.6 41.3 72.3 66.6 70.4
CoDA (Ours) 71.8 60.9 61 66.4 61.2 59.2 41.9 75.3 70.9 72.6
@@ 4.2 State-of-the-art Performance Comparison
Cityscapes to rainy, snowy, and all scenes We conduct CS to ACDC-All and test on ACDC-Rain, Snow, and All for these experiments. As shown in Table 1, CoDA achieves 72.6% mIoU on the ACDC-All benchmark outperforming the MIC 2.2%, and demonstrates strong generalizability on ACDC-Rain and Snow with improvements of 3.0% and 4.3% respectively compared to MIC. For more details, we show the results of every class of CS to ACDC-All in Table 2. As illustrated in the Intro section, the previous methods show weakness in recognizing the sky under night scenes, which is alleviated by our CoDA with 8.4% improvements in the sky class compared to MIC.
@@ 4.3 Ablation Study Analysis
Although the adding position of visual prompts is discussed in [2, 12, 58], the discussion of the initial parameters of visual prompts is relatively blank. Thus, we conduct joint experiments of positions with different initializations shown in Table 5. During 40$\sim$60k, almost each experiment improves compared to their 0$\sim$40k scores and the Corner+Center shows the best performance with 72.1% mIoU. When initial $\delta_{v}$=1, Padding Patch and Corner+Center performances reduce a lot contrary to Random Patch with 70.8% mIoU. We argue that $\delta_{v}$=1 gives models a noisy begin and enough randomness can alleviate the noise influences. Also, the initialization with uniform and normal distribution achieves similar mIoU, which means MIC is not sensitive to initial parameter distributions and results of $\delta_{v}$=0 demonstrate that starting from scratch is the best choice for Meta-Visual Prompts. The visualization details of three positions with different initializations are attached in the supplementary material.
@@
Ablation studies of CoDA components We conduct experiments to show the necessity of each component in CoDA in Table 8. we combine 1, 2, and 4 rows to reveal that the prerequisite of SAVPT is the good foundation brought by the CoD+traditional strategy. From 4, 5, and 7 rows, we know that $X_{G}$ and SAVPT need to collaborate to maximize the effects with 0.1% to 3.8% gain from 4 to 7 rows and 2.1% to 3.8% gain from 5 to 7 rows. Combining 6 and 7 rows, results show that SPT is decisive for CoDA’s performance bringing 4.2% improvements from 67.5 to 72.1 mIoU. Further details about CoDA’s performance, ablation studies, and motivations are attached in our supplementary material.
@@ 5 Conclusion
In this paper, we propose a method named CoDA consists of a CoD strategy with a tailored dataset and SAVPT mechanism. Extensive experiments validate that CoD provides scene-level instructions to models for eliminating hallucinations on tough scenes, while SAVPT serves as a plug-in mechanism enhancing models’ inherent abilities without complicating networks through image-level instructions. In this paper, however, we do not detailedly discuss why the Meta-adapter emerges attributes akin to visual prompts that can be discarded during training. This important phenomenon will be one of our future investigations.
@@ 0.B.2 The Generated Dataset
Visual representations of images depicting fog, snow, and rain, accompanied by the corresponding four stages of phrasal prompts $\Phi$, are depicted in Figures 7, 8, and 9, respectively. In the initial stage of our experiments, we generate $\Phi_{1}$ randomly for each scenario by employing phrasal weather prompts, such as "<rainy/snowy/foggy> street view," alongside four randomly chosen prompts from $\Phi_{1}$ across all scenes. For subsequent stages of generation, we adhere to the progression from $\Phi_{2}$ through $\Phi_{4}$, with edits limited only to the weather conditions and number of objects depicted in the prompts. Upon integrating $\Phi_{1}$ and $\Phi_{2}$, it becomes evident that images produced using $\Phi_{2}$ generally demonstrate cleaner and better texture details compared to those generated from $\Phi_{1}$. Specifically, the sky generated in images from $\Phi_{2}$ surpasses the quality of those from random $\Phi_{1}$, notably avoiding the erroneous inclusion of traffic lights in the sky. Furthermore, the $\Phi_{2}$ generated images are characterized by a reduction in anomalous objects, such as low-quality depictions of cyclists and trains misplaced on roads, when compared to $\Phi_{1}$. These findings suggest that $\Phi_{2}$ prompts, which encapsulate the shared and basic features of both source and target domains, are more beneficial to the performance of Stable Diffusion 2 (SD2) than the random $\Phi_{1}$ prompts that focus solely on target domain features. This evidence validates the effectiveness of the second stage in our methodology and aligns well with our above motivation that the common and basic features indeed can strengthen the robustness of SD2.

### v3

@@ 2.1 Semantic segmentation under adverse scenes within UDA
Since existing works [57, 58, 69, 60, 61] predominately focus on four adverse scenes (fog, rain, snow, and night), we divide them into three branches by the scene benchmarks that they evaluate. The first branch comprises methods focusing on parsing foggy scenes [43, 44, 9] trying to generate synthetic fog images to narrow the real-to-foggy gaps. Later works tried other paths like CuDA-Net [36] that reported disentangling gaps by introducing an intermediate domain, and FIFO [31] that presented an auxiliary network to help the segmentation model to learn fog-invariant features. Since the above works generalize well to rainy and snowy scenes, no methods are specializing in rainy and snowy scenes. Night-specialized methods [59] are contained in the second branch and they too focus on narrowing the gaps by introducing extra domain images like twilight images or extra information[10, 42, 45]. The final branch comprises models trying to solve multiple scenes including normal and adverse scenes. DAFormer [23] was the first to introduce the SegFormer-based [63] architecture to this field. Subsequently, some DAFormer-based methods [62, 4] emerged like STA [19] which achieves domain generalization in UDA, and HRDA [24] that proposes multi-scale features fusion network. Some works also take HRDA as the baseline model, such as MIC [25] using masks to help models learn context-level features. Different from previous methods, CoDA is the first UDA method to focus models on differences within adverse scenes to learn domain-invariant features.
2.2 CoT and its variants
@@
2.3 Visual prompt tuning in UDA
Prompt-based learning [34] initially serving as Parameter-Efficient Fine-Tuning (PEFT) [7, 68, 22] is used to finetune large pre-trained models to downstream tasks in Natural Language Processing (NLP). Subsequently, the prompt has begun to be transferred to the field of Computer Vision (CV) bringing about a significant academic trend. Many methods [38, 29, 65, 70] first tried to migrate prompt to Vision Language Models (VLM) in text form, and then VPT [28] first introduced "visual prompt" into CV as learnable vectors (soft prompt). [17] followed by other works [16, 15, 47] is the first work focusing on domain adaptation and employing prompt. Most relevant to our work are the methods that focus on UDA or DG for adverse scene understanding, such as [14] employs text features to strengthen features achieving a novel style augmentation implicitly, and [51] combines clip [38] to conduct semantic augmentation.
@@ 3.3 Instructive Chain-of-Domain Adaptation
The first key point in scene-level CoD aims to simply divide all adverse scenes into Easy and Hard two-level scenes to construct two intermediate domains, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Take Cityscapes to ACDC experiment as an example, ACDC representing the target domain $T$ consists of fog, rain, snow, and night scenes, so we directly choose daytime scene images (fog, rain, snow, and $X_{G}$) as easy scene images to construct $X_{\mathcal{M}_{1}}$ and use night images to build $X_{\mathcal{M}_{2}}$. The data composition of Cityscapes to ACDC can be represented as Fig 3 (a). Thus, the CoD adaptation process transforms from $S\rightarrow T$ into $S\rightarrow\mathcal{M}_{1}\rightarrow\mathcal{M}_{2}\rightarrow T^{\prime}$.
@@ 4.1 Datasets and Benchmarks
Cityscapes (CS) [8] is a real-world dataset captured with normal driving scenes in 50 urban areas. Foggy Zurich (FZ) [43] and Foggy Driving (FD) [44] are two foggy real-world datasets. FZ contains 3808 images with light and medium fog. Notably, only 40 images are for testing in FZ. Unlike FZ, FD is a dataset purely for testing as it contains 101 labeled images. Dark Zurich (DZ) [42] is captured during nighttime, twilight, and daytime including 50 validation images and 151 test images. Nighttime Driving (ND) [10] contains 50 nighttime images with coarsely annotated ground truth. BDD100K-Night (BD) [66] is a subset of the BDD100K, we only use 87 test images. ND is also used for testing. ACDC [46] is a dataset under fog, rain, snow, and night scenes. There are 400 training images, 100 validation images (106 in night), and 500 testing images in every scene. Besides, ACDC also contains 1600 clean images called ACDC-ref.
@@
Models Backbone Adverse Scenes (mIoU)
@@
SFSU [44] RefineNet [33] - 35.7 46.3 - - - - - - -
CMAda3 [9] RefineNet - 46.8 49.8 - - - - - - -
CuDA-Net [36] DeepLab-v2 [5] 55.6 49.1 53.5 - - - - - - -
FIFO [31] RefineNet - 48.4 50.7 - - - - - - -
FogAdapt [26] ResNet-38 - 50.6 53.4 - - - - - - -
SAM-EDA [55] ViT-H [12] - - 56.4 - - - - - - -
GCMA [42] RefineNet - - - - 42.0 45.6 33.2 - - -
MCGDA [45] RefineNet - - - - 42.5 49.4 34.9 - - -
DANNet [59] RefineNet - - - - 44.3 47.7 - - - 50.0
@@
AdaptSeg [49] DeepLab-v2 - 26.1 37.6 - 30.4 34.5 22.0 - - -
DAFormer [23] SegFormer [63] 48.9 40.8 - 44.7 48.5 51.8 33.9 59.9 53.7 55.4
SePiCo [62] SegFormer 58.5 - - 50.5 54.2 56.9 40.6 66.1 57.9 59.1
STA [19] SegFormer 60.2 46.9 54.9 48.4 - - - 61.3 58.0 60.9
HRDA [24] SegFormer 69.9 46.0 - 53.1 55.9 - - 73.6 69.5 68.0
MIC [25](baseline) SegFormer 67.0 53.3 56.6 57.2 60.2 58.6 41.3 72.3 66.6 70.4
CoDA (Ours) SegFormer 71.8 60.9 61 66.4 61.2 59.2 41.9 75.3 70.9 72.6
@@ 4.2 State-of-the-art Performance Comparison
Cityscapes to rainy, snowy, and all scenes We conduct CS to ACDC-All and test on ACDC-Rain, Snow, and All for these experiments. As shown in Table 1, CoDA achieves 72.6% mIoU on the ACDC-All benchmark outperforming the MIC 2.2%, and demonstrates strong generalizability on ACDC-Rain and Snow with improvements of 3.0% and 4.3% respectively compared to MIC. For more details, we show the results of every class of CS to ACDC-All in Table 2. As illustrated in the Intro section, the previous methods show weakness in recognizing the sky under night scenes, which our CoDA alleviates with 8.4% improvements in the sky class compared to MIC.
@@ 4.3 Ablation Study Analysis
Although the adding position of visual prompts is discussed in [2, 15, 67], the discussion of the initial parameters of visual prompts is relatively blank. Thus, we conduct joint experiments of positions with different initializations shown in Table 4. During 40$\sim$60k, almost each experiment improves compared to their 0$\sim$40k scores and the Corner+Center shows the best performance with 72.1% mIoU. When initial $\delta_{v}$=1, Padding Patch and Corner+Center performances reduce a lot contrary to Random Patch with 70.8% mIoU. We argue that $\delta_{v}$=1 gives models a noisy begin and enough randomness can alleviate the noise influences. Also, the initialization with uniform and normal distribution achieves similar mIoU, which means MIC is not sensitive to initial parameter distributions and results of $\delta_{v}$=0 demonstrate that starting from scratch is the best choice for Meta-Visual Prompts.
@@ 5 Conclusion
In this paper, we propose a method named CoDA, which consists of a CoD strategy with a tailored dataset and SAVPT mechanism. Extensive experiments validate that CoD provides scene-level instructions to models for eliminating hallucinations on tough scenes, while SAVPT serves as a plug-in mechanism enhancing models’ inherent abilities without complicating networks through image-level instructions. In this paper, however, we do not detailedly discuss why the Meta-adapter emerges attributes akin to visual prompts that can be discarded during training. This important phenomenon will be one of our future investigations.
Acknowledgement
This study was funded by the National Natural Science Foundation of China (Grants No. U21A6001) and the Innovation Group Project of Southern Marine Science and Engineering Guangdong Laboratory (Zhuhai) (Grants No.     SML2023SP208). We also acknowledge the high-performance computing support from School of Atmospheric Science at Sun Yat-sen University.
@@ Appendix 0.A Additional Results
Num MIC CoD+Tra. $X_{G}$ VPT SAVPT mIoU Gain
1 - - - - 68.3 -
2 - - - 64.3 $-4.0$
3 - - - 70.3 $+2.0$
4 - - 68.4 $+0.1$
5 - - 70.4 $+2.1$
6 - 67.5 $-0.8$
7 - 72.1 $+3.8$
Ablation studies of CoDA components We conduct experiments to show the necessity of each component in CoDA in Table 8. we combine 1, 2, and 4 rows to reveal that the prerequisite of SAVPT is the good foundation brought by the CoD+traditional strategy. From 4, 5, and 7 rows, we know that $X_{G}$ and SAVPT need to collaborate to maximize the effects with 0.1% to 3.8% gain from 4 to 7 rows and 2.1% to 3.8% gain from 5 to 7 rows. Combining 6 and 7 rows, results show that SPT is decisive for CoDA’s performance bringing 4.2% improvements from 67.5 to 72.1 mIoU.
@@ 0.B.2 The Generated Dataset
Visual representations of images depicting fog, snow, and rain, accompanied by the corresponding four stages of phrasal prompts $\Phi$, are depicted in Figures 7, 8, and 9, respectively. In the initial stage of our experiments, we generate $\Phi_{1}$ randomly for each scenario by employing phrasal weather prompts, such as "<rainy/snowy/foggy> street view," alongside four randomly chosen prompts from $\Phi_{1}$ across all scenes. For subsequent stages of generation, we adhere to the progression from $\Phi_{2}$ through $\Phi_{4}$, with edits limited only to the weather conditions depicted in the prompts. Upon integrating $\Phi_{1}$ and $\Phi_{2}$, it becomes evident that images produced using $\Phi_{2}$ generally demonstrate cleaner and better texture details compared to those generated from $\Phi_{1}$. Specifically, the sky generated in images from $\Phi_{2}$ surpasses the quality of those from random $\Phi_{1}$, notably avoiding the erroneous inclusion of traffic lights in the sky. Furthermore, the $\Phi_{2}$ generated images are characterized by a reduction in anomalous objects, such as low-quality depictions of cyclists and trains misplaced on roads, when compared to $\Phi_{1}$. These findings suggest that $\Phi_{2}$ prompts, which encapsulate the shared and basic features of both source and target domains, are more beneficial to the performance of Stable Diffusion 2 (SD2) than the random $\Phi_{1}$ prompts that focus solely on target domain features. This evidence validates the effectiveness of the second stage in our methodology and aligns well with our above motivation that the common and basic features indeed can strengthen the robustness of SD2.